OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 4 points) [9.01] When the quadratic equation y = (x − 4)(2x + 5) is written in standard form, what is the value of the coefficient "b"? 2 3 −3 −20 Question 2 (Multiple Choice Worth 4 points) [9.01] Identify which equation is a quadratic equation. y = −6x − 3 y + x2 = (x − 1)(x + 7) y − 9 = (x − 1)(x2 + 4) y = 5x(x + 3) − 8 Question 3 (Multiple Choice Worth 4 points) [9.02] What are the solutions of x2 − 3x − 18 = 0? x = 3, x = −6 x = −3, x = 6 x = 3, x = 6 x = −3, x = −6

OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 4 points) [9.01] When the quadratic equation y = (x − 4)(2x + 5) is written in standard form, what is the value of the coefficient "b"? 2 3 −3 −20 Question 2 (Multiple Choice Worth 4 points) [9.01] Identify which equation is a quadratic equation. y = −6x − 3 y + x2 = (x − 1)(x + 7) y − 9 = (x − 1)(x2 + 4) y = 5x(x + 3) − 8 Question 3 (Multiple Choice Worth 4 points) [9.02] What are the solutions of x2 − 3x − 18 = 0? x = 3, x = −6 x = −3, x = 6 x = 3, x = 6 x = −3, x = −6 Question 4 (Multiple Choice Worth 4 points) [9.02] What are the x-intercept(s) of the graph of y − x2 = −8x − 9? (−1, 0) and (9, 0) (1, 0) and (−9, 0) (−1, 0) and (−9, 0) (1, 0) and (9, 0) Question 5 (Multiple Choice Worth 4 points) [9.02] What are the solutions of 6x2 − x − 2 = 0? x = , x = − x = −, x = x = , x = −1 x = −, x = 2 Question 6 (Multiple Choice Worth 4 points) [9.02] Solve 2x2 = −10x. x = 2, x = −5 x = 2, x = 5 x = 0, x = 5 x = 0, x = −5 Question 7 (Multiple Choice Worth 4 points) [9.01] What is the vertex for the graph of y = 3x2 − 12x + 9? (−2, 21) (−2, 45) (2, −3) (2, −9) Question 8 (Multiple Choice Worth 4 points) [9.01] Determine whether the graph of y = 2x2 − 5x + 8 opens up or down and whether it has a maximum or minimum point. Down and Maximum Down and Minimum Up and Minimum Up and Maximum Question 9 (Multiple Choice Worth 4 points) [9.01] The graph of which equation is shown below? http://learn.flvs.net/webdav/assessment_images/educator_algebra1_v10/09_04_17.jpg y = x2 − 4x + 1 y = x2 + 4x + 1 y = −x2 − 4x + 1 y = −x2 + 4x + 1 Question 10 (Essay Worth 7 points) [9.01], [9.02] Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x2 – 36 = 0 and x2 = 8x – 12 Part 2: Describe what the solution(s) represent to the graph of each. (2 points) Part 3: How are the graphs alike? How are they different? (2 points) Question 11 (Essay Worth 7 points) [9.01], [9.02] Create your own unique quadratic equation in the form y = ax2 + bx + c. Use complete sentences and show all work to determine the following: Does the graph open up or down? How do you know? (2 points) Explain whether the graph has a maximum or minimum point. (2 points) Find the vertex and x-intercepts of the graph. (3 points)

OpenStudy (anonymous):

wooooooooooow .-.

OpenStudy (anonymous):

I suck at math. big time. lol

OpenStudy (anonymous):

i hate math

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

just google it

OpenStudy (anonymous):

that didnt help, and google doesnt help either... -.-

OpenStudy (anonymous):

why how when where who